OpenStudy (anonymous):

A T shirt launcher launches shirts at 15 m/s. If it takes 0.30 s to accelerate a 250 g shirt to launch speed, determine the force exerted on the shirt. You can assume no friction. I just need the final answer Please !

OpenStudy (anonymous):

a = f.v - i.v /t a = 15 - 0 / 0.30 a 50ms^-2 F = ma...Let R represent frictional force F - R = ma F - 0 = 0.250kg * 50ms^-2 Find the force :)